# Generalized Lotka-Volterra model with sparse interactions: non-Gaussian effects and topological multiple-equilibria phase

### v1

@@ V Conclusions
In this work we have studied the equilibrium thermodynamics of the generalized Lotka-Volterra (gLV) model on a sparse random graph with quenched disordered interactions, with particular attention to the behaviour of species abundance distributions. By exploiting the cavity method and the Belief Propagation equations, we have studied the properties of the gLV model on a random regular graph with small and fixed connectivity $k$, varying three parameters: the average interaction strength $\hat{\mu}$, the width $\hat{\sigma}$ of the randomly distributed disordered couplings and the extent of thermal fluctuations, parametrized by the temperature $T$. Our first important observation has been that, by increasing $\hat{\sigma}$ at finite and small $\hat{\mu}$ and finite temperature, the abundance marginal distributions become strongly non-Gaussian while staying in a single-equilibrium phase. In models with the same sort of random couplings but living on a dense network, there is by construction no room for non-Gaussian effects. We have thus argued that this result is purely due to the sparsity of the network, therefore providing a new perspective on the emergence of non-Gaussian and Gamma-like distributions in ecological systems. In particular, the finding of Gamma-like marginals is strongly consistent with the empirical observations from real ecosystems [13, 14]. We have observed non-Gaussian effects also in two-species joint probabilities, due to the emergence of non trivial correlations. In addition to the observation of these non-Gaussian effects in the single equilibrium phase, we have ascertained that by further increasing $\hat{\sigma}$ for the same choice of $\hat{\mu}$ and $T$ the glass phase encountered in fully-connected networks is never met: on the Bethe lattice with fixed and small connectivity, at small values of $\hat{\mu}$, upon increasing $\hat{\sigma}$ the system has a transition from the single-equilibrium phase directly to the unbounded growth one.
@@
In the second part of our study we concentrated on the phenomenology of the gLV model on a sparse topology and with zero disorder, i.e., $\hat{\sigma}=0$. By varying both the temperature $T$ and the average strength $\hat{\mu}$ of the ordered interactions we have been thus able to extend the zero temperature results of [17], showing that the formation of the topological glass phase pointed out in [17] at $T=0$ for high enough interaction strength, $\hat{\mu}>\hat{\mu}_{c}$, is robust against the introduction of temperature. This phase emerges due to the occurrence of extinctions, resulting from high competition within the system (large $\hat{\mu}$). The higher is $\hat{\mu}$, the larger is the amount of extinctions, which eventually determines the formation of disconnected groups of surviving species and the freezing of the system into a configuration where a certain number of species have gone extinct, forming non-communicating islands of surviving ones. The pattern of surviving/disappearing species strongly depends on the initial conditions and the sequence of thermal fluctuations, so that there is a huge multiplicity of possible target stationary states: that is why it makes sense to talk about a glass phase, but of very different nature with respect to the one induced by the disordered couplings in fully connected networks [2].
@@
In conclusion, the main outcomes of our investigation have been the highlight of realistic non-Gaussian effects in the single-equilibrium phase and the uncovering of a non-trivial glass phase at finite temperature and in the absence of disorder, therefore questioning two of the main ingredients of the models so-far used to mimic the behaviour of large ecosystems: dense networks and large fluctuations in the disordered couplings. From a model-building perspective, we provided the evidence that the study of sparse networks offers a new avenue for understanding the behaviour of large ecosystems within a more realistic framework.
@@ Acknowledgments
We thank David Machado Perez and Pietro Valigi for useful discussions. We acknowledge support from the project MIUR-PRIN2022, “Emergent Dynamical Patterns of Disordered Systems with Applications to Natural Communities”, code 2022WPHMXK_003, funded by European Union — Next Generation EU, Mission 4 Component 1 CUP:D53D23002900006.
@@ Appendix A Cavity Equations and Belief Propagation Algorithm
where $z_{i\rightarrow j}$ and $z_{i}$ are normalization factors that ensure $\sum_{n_{i}}\eta_{i\rightarrow j}(n_{i})=\sum_{n_{i}}\eta_{i}(n_{i})=1$. One can easily show that Eqs. (A) are exact on a tree exploiting its structure. Indeed, on a tree, for each node $i$, the messages $\eta_{k\rightarrow i}(n_{k})$ coming from the neighbours $k$ of $i$ are all independent from each other once the edge $(i,k)$ has been cut, and so their contributions can be simply multiplied. This is not necessarily true in the case of RRG’s. In particular in such graphs loops are present and so the messages $\eta_{k\rightarrow i}(n_{k})$ of the neighbours $k$ of node $i$ could be in principle not independent from each other when we cut the edges $(i,k)$. However, in RRG’s the typical length of loops is $O(\log N)$ and thus RRG are locally tree like (Fig.11) and in the thermodynamic limit Eqs (A) are exact also for RRG’s, as long as a the whole Gibbs measure is composed of just one state. In particular for $N$ large enough we can use Eqs. (A) also for our gLV-system on a sparse random regular graph.
@@
$\eta_{i\rightarrow j}(n_{i})=\frac{1}{z_{i\rightarrow j}}\exp{\left(\beta r\left(n_{i}-\frac{n_{i}^{2}}{2K}\right)-\ln\,(n_{i}+\epsilon)\right)}\prod_{k\in\partial i\backslash j}\left(\sum_{\{n_{k}\}}\eta_{k\rightarrow i}(n_{k})\exp{-\beta\alpha_{ik}\frac{r}{K}n_{i}n_{k}}\right),$ ) ) , (28)
@@
$\eta_{i}(n_{i})=\frac{1}{z_{i}}\exp{\left(\beta r\left(n_{i}-\frac{n_{i}^{2}}{2K}\right)-\ln\,(n_{i}+\epsilon)\right)}\prod_{j\in\partial i}\left(\sum_{\{n_{j}\}}\eta_{j\rightarrow i}(n_{j})\exp{-\beta\alpha_{ij}\frac{r}{K}n_{i}n_{j}}\right).$ ) ) . (29)
The marginal distribution $\eta_{i}(n_{i})$ thus represents the probability for a species $i$ to have a number $n_{i}$ of individuals, once we have marginalized over all the other species. Once we have the set of equations for the cavity marginals given by (28), we can solve them iteratively. In principle, the cavity marginals could be thought as continuous probability distributions. However, to practically solve the self-consistent equations, it is much more useful to discretize the values that can be taken by the variable $n_{i}$. In particular, we will assume that $n_{i}\in[0,1,...,n_{\text{max}}-1]$, and we will take $n_{\text{max}}$ large enough so that $\forall i\quad\eta_{i}(n_{\text{max}}-1)=0$ inside the numerical precision of a computer. If for any value of $n_{\text{max}}$, the value $\eta_{i}(n_{\text{max}})$ is larger than zero (inside computer precision), then the species $i$ is considered to be in the unbounded growth phase. For each oriented edge $i\rightarrow j$, the cavity marginal $\eta_{i\rightarrow j}(n_{i})$ has to be initialized to a normalized distribution. For example we can take, for every oriented couple $(i,j)$, $\eta_{i\rightarrow j}^{(0)}(n_{i})=\frac{1}{n_{\text{max}}}$, $\forall\,n_{i}\in[0,1,...,n_{\text{max}}-1]$, i.e. an uniform initial condition. But in general, we checked that every result of the paper is independent on the chosen initial conditions. Starting from $\eta_{i\rightarrow j}^{(0)}$, we can compute at the following time step the normalization factor $z_{i\rightarrow j}^{(1)}$ and the cavity marginals $\eta_{i\rightarrow j}^{(1)}$. In general, from step $t$ to step $t+1$ the iteration process goes like this:
$z_{i\rightarrow j}^{(t+1)}$ =$=\sum_{\{n_{i}\}}\exp{\left(\beta r\left(n_{i}-\frac{n_{i}^{2}}{2K}\right)-\ln\,(n_{i}+\epsilon)\right)}\prod_{k\in\partial i\backslash j}\left(\sum_{\{n_{k}\}}\eta_{k\rightarrow i}^{(t)}(n_{k})\exp{-\beta\alpha_{ik}\frac{r}{K}n_{i}n_{k}}\right),$ ) ) , (30)
$\eta_{i\rightarrow j}^{(t+1)}(n_{i})$ $=\frac{1}{z_{i\rightarrow j}^{(t+1)}}\exp{\left(\beta r\left(n_{i}-\frac{n_{i}^{2}}{2K}\right)-\ln\,(n_{i}+\epsilon)\right)}\prod_{k\in\partial i\backslash j}\left(\sum_{\{n_{k}\}}\eta_{k\rightarrow i}^{(t)}(n_{k})\exp{-\beta\alpha_{ik}\frac{r}{K}n_{i}n_{k}}\right).$ ) ) . (31)
@@
Figure 12: Single species marginal distributions $\eta_{i}(n)$ for two different values of $\hat{\sigma}$, with $\hat{\mu}=0.1$, $T=1$ and $N=256$. In the Gaussian regime ($\hat{\sigma}=0.02$) the single marginals are Gaussian and concentrated within a small range of $n$. In the non-Gaussian regime ($\hat{\sigma}=0.20$) the distributions spread over a larger range of $n$, extinctions appear, and not all single marginals are Gaussian anymore.
Appendix B Universality and Scaling Law for the Sparse Ecosystems Dynamics

### v2

@@ V Conclusions
In this work we have studied the equilibrium thermodynamics of the generalized Lotka-Volterra (gLV) model on a sparse random graph with quenched disordered interactions, with particular attention to the behaviour of species abundance distributions. By exploiting the cavity method and the Belief Propagation equations, we have studied the properties of the gLV model on a random regular graph with small and fixed connectivity $k$, varying three parameters: the average interaction strength $\hat{\mu}$, the width $\hat{\sigma}$ of the randomly distributed disordered couplings and the extent of thermal fluctuations, parametrized by the temperature $T$. Our first important observation has been that, by increasing $\hat{\sigma}$ at finite and small $\hat{\mu}$ and finite temperature, the abundance marginal distributions become strongly non-Gaussian while staying in a single-equilibrium phase. In models with the same sort of random couplings but living on a dense network, there is by construction no room for non-Gaussian effects. We have thus argued that this result is purely due to the sparsity of the network, therefore providing a new perspective on the emergence of non-Gaussian and Gamma-like distributions in ecological systems. In particular, the finding of Gamma-like marginals is strongly consistent with the empirical observations from real ecosystems [13, 19]. We have observed non-Gaussian effects also in two-species joint probabilities, due to the emergence of non trivial correlations. In addition to the observation of these non-Gaussian effects in the single equilibrium phase, we have ascertained that by further increasing $\hat{\sigma}$ for the same choice of $\hat{\mu}$ and $T$ the multiple attractor phase encountered in fully-connected networks is never met: on the Bethe lattice with fixed and small connectivity, at small values of $\hat{\mu}$, upon increasing $\hat{\sigma}$ the system has a transition from the single-equilibrium phase directly to the unbounded growth one.
@@
In the second part of our study we concentrated on the phenomenology of the gLV model on a sparse topology and with zero disorder, i.e., $\hat{\sigma}=0$. By varying both the temperature $T$ and the average strength $\hat{\mu}$ of the ordered interactions we have been thus able to extend the zero temperature results of [16], showing that the formation of the topological multiple equilibria phase pointed out in [16] at $T=0$ for high enough interaction strength, $\hat{\mu}>\hat{\mu}_{c}$, is robust against the introduction of temperature. This phase emerges due to the occurrence of extinctions, resulting from high competition within the system (large $\hat{\mu}$). The higher is $\hat{\mu}$, the larger is the amount of extinctions, which eventually determines the formation of disconnected groups of surviving species and the freezing of the system into a configuration where a certain number of species have gone extinct, forming non-communicating islands of surviving ones. The pattern of surviving/disappearing species strongly depends on the initial conditions and the sequence of thermal fluctuations, so that there is a huge multiplicity of possible target stationary states: that is why it makes sense to talk about a multiple attractor phase, but of very different nature with respect to the one induced by the disordered couplings in fully connected networks [1].
@@
In conclusion, the main outcomes of our investigation have been the highlight of realistic non-Gaussian effects in the single-equilibrium phase and the uncovering of a non-trivial multiple attractor phase at finite temperature and in the absence of disorder, therefore questioning two of the main ingredients of the models so-far used to mimic the behaviour of large ecosystems: dense networks and large fluctuations in the disordered couplings. From a model-building perspective, we provided the evidence that the study of sparse networks offers a new avenue for understanding the behaviour of large ecosystems within a more realistic framework.
@@ Acknowledgments
We thank David Machado Perez and Pietro Valigi for useful discussions. We acknowledge support from the project MIUR-PRIN2022, “Emergent Dynamical Patterns of Disordered Systems with Applications to Natural Communities”, code 2022WPHMXK, funded by European Union — Next Generation EU, Mission 4 Component 1 CUP: C53D23001560006, D53D23002900006 and B53D23005310001.
@@ Appendix A Cavity Equations and Belief Propagation Algorithm
where $z_{i\rightarrow j}$ and $z_{i}$ are normalization factors that ensure $\sum_{n_{i}}\eta_{i\rightarrow j}(n_{i})=\sum_{n_{i}}\eta_{i}(n_{i})=1$. One can easily show that Eqs. (A) are exact on a tree exploiting its structure. Indeed, on a tree, for each node $i$, the messages $\eta_{k\rightarrow i}(n_{k})$ coming from the neighbours $k$ of $i$ are all independent from each other once the edge $(i,k)$ has been cut, and so their contributions can be simply multiplied. This is not necessarily true in the case of RRG’s. In particular in such graphs loops are present and so the messages $\eta_{k\rightarrow i}(n_{k})$ of the neighbours $k$ of node $i$ could be in principle not independent from each other when we cut the edges $(i,k)$. However, in RRG’s the typical length of loops is $O(\log N)$ and thus RRG are locally tree like (Fig.12) and in the thermodynamic limit Eqs (A) are exact also for RRG’s, as long as a the whole Gibbs measure is composed of just one state [20]. In particular for $N$ large enough we can use Eqs. (A) also for our gLV-system on a sparse random regular graph.
@@
[…]
@@
[…]
The marginal distribution $\eta_{i}(n_{i})$ thus represents the probability for a species $i$ to have a number $n_{i}$ of individuals, once we have marginalized over all the other species. Once we have the set of equations for the cavity marginals given by (28), we can solve them iteratively. In principle, the cavity marginals could be thought as continuous probability distributions. However, to practically solve the self-consistent equations, it is much more useful to discretize the values that can be taken by the variable $n_{i}$. In particular, we will assume that $n_{i}\in[0,1,...,n_{\text{max}}-1]$, and we will take $n_{\text{max}}$ large enough so that $\forall i\quad\eta_{i}(n_{\text{max}}-1)=0$ inside the numerical precision of a computer. For each oriented edge $i\rightarrow j$, the cavity marginal $\eta_{i\rightarrow j}(n_{i})$ has to be initialized to a normalized distribution. For example we can take, for every oriented couple $(i,j)$, $\eta_{i\rightarrow j}^{(0)}(n_{i})=\frac{1}{n_{\text{max}}}$, $\forall\,n_{i}\in[0,1,...,n_{\text{max}}-1]$, i.e. an uniform initial condition. But in general, we checked that every result of the paper is independent on the chosen initial conditions. Starting from $\eta_{i\rightarrow j}^{(0)}$, we can compute at the following time step the normalization factor $z_{i\rightarrow j}^{(1)}$ and the cavity marginals $\eta_{i\rightarrow j}^{(1)}$. In general, from step $t$ to step $t+1$ the iteration process goes like this:
[…]
@@
Figure 13: Single species marginal distributions $\eta_{i}(n)$ in the Gaussian regime (left) and in the non-Gaussian regime (right). In the Gaussian regime ($\hat{\sigma}=0.02$) the single marginals are Gaussian and concentrated within a small range of $n$. In the non-Gaussian regime ($\hat{\sigma}=0.20$) the distributions spread over a larger range of $n$, extinctions appear, and not all single marginals are Gaussian anymore. The other parameters are $\hat{\mu}=0.1$, $T=1$ and $N=256$.
Appendix B Independence of the marginals from the regularization parameter $\epsilon$ and the discretization step
As explained in Appendix A, in our approach, in order to numerically solve the BP equations we chose to discretize the abundance distributions into nonnegative integers. We made this choice because in this way one can interpret $n_{i}$ as the number of individuals for the species $i$ (that is, by definition, an integer). However, this is not an inevitable choice: indeed one can choose to discretize $n_{i}$ into small steps. We do not expect that the integer discretization influences the results because indeed the discretization step, that in our case is $dn=1$ should be compared with the natural scale for $n$ that is $K$ (fixed point of the single species dynamics) that we chose to be large exactly for this reason ($K=280$). To verify this, in this Appendix we show that the results do not change changing $dn$. In particular, in Fig. 14 we show the average marginal distribution with the same two set of parameters used in Fig. 1 of the main text, but with a discretization step $dn=0.5$, that gives non integer numbers for the abundances. The distributions are completely equivalent to the ones with $dn=1$, for the exception of a very small variation at small values of $n$, that leads to a marginal distribution with $dn=0.5$ that is more comparable to the results of the dynamics than the one with $dn=1$ at small $n$.
Figure 14: Comparison between species abundances computed with discretization steps $dn=1$ and $dn=0.5$. For small $\hat{\sigma}=0.02$ (blue and black points) the distributions with $dn=1$ and $dn=0.5$ are perfectly equivalent, for large $\hat{\sigma}=0.20$ (orange and red points) there are small differences between the two distributions at small values of $n$. Inset: marginal at $\hat{\sigma}=0.20$ computed using BP with $dn=0.5$ (red points) and from the dynamics (green points). The two agree very well and the BP marginal computed with $dn=0.5$ is more similar to the dynamics than the one computed with $dn=1$. In both plots the parameters are the same as in Fig. 1 of the main text: $T=1$, $\hat{\mu}=0.1$, $N=256$.
Figure 15: Left: Comparison between species abundance distributions computed on a single graph with different regularization parameters $\epsilon$. For small $\hat{\sigma}=0.02$ the distributions with $\epsilon=10^{-4},10^{-5},10^{-6}$ are perfectly equivalent, for large $\hat{\sigma}=0.20$ there are small differences between the three distributions at small values of $n$. Inset: Same as in the main Figure, but in logarithmic scale to highlight the small differences at small $n$. Right: Surviving probability $\Phi$ as a function of $\hat{\sigma}$ for different values of $\epsilon$. In both plots the parameters are: $T=0.5$, $\hat{\mu}=0.1$, $N=128$.
As explained in the main text, because of the discretized version of BP, we replaced the reflecting wall used for example in ref. [1], introducing a regularization parameter $\epsilon$ in the term $T\ln(n_{i})$ inside the Hamiltonian of eq. (II), otherwise ill-defined at $n_{i}=0$. We show now that the obtained results for the marginal distributions are almost independent on the choice of $\epsilon$. In the left panel of Fig. 15 we show the average species abundance for two different values of the variance of the couplings $\hat{\sigma}$ for three different values of the parameter $\epsilon$. For small $\hat{\sigma}$ the distributions with $\epsilon=10^{-4},10^{-5},10^{-6}$ are perfectly equivalent, while for large $\hat{\sigma}=0.20$ there are very small differences between the three distributions at small values of $n$ (only visible in log scale).
In the right panel of Fig. 15 we show how the Surviving probability $\Phi$, defined as $\Phi=1-\eta(0)$, changes for different values of $\epsilon$. The point at which $\Phi$ becomes different from 1 is independent of $\epsilon$, while for larger $\hat{\sigma}$, where extinctions are possible, $\Phi$ varies of a few percent passing from $\epsilon=10^{-4}$ to $\epsilon=10^{-6}$. However, we verified that the transition from the single fixed point phase to the unbounded growth phase (at small values of $\hat{\mu}$, such the one in Fig. 15) is independent on $\epsilon$. Moreover, in Appendix G, we will show that also the transition from the single fixed point phase to the multiple attractor phase at high values of $\hat{\mu}$ is independent on the value of $\epsilon$.
Appendix C Detecting the Unbounded Growth phase
As said in the previous section, in order to have a numerical implementation of the BP equations, we both discretize $n$ and introduce a maximum value $n_{max}$. In this section we show how the final marginals are independent on $n_{max}$ (for $n_{max}$ sufficiently high) in the single fixed point phase, and we show how to detect the Unbounded Growth phase.
Figure 16: Left: Marginals $\mathbf{\eta_{i}^{\alpha}(n)}$ for three species changing $n_{max}$ in the unique fixed point phase. In the unique fixed point phase, if $n_{max}$ is too small, the marginals for some species will be wrong: We detect the error looking if there are some species for which $\eta_{i}(n_{max}-1)\neq 0$, as happens for species 1 in the figure. Enlarging $n_{max}$ will then correct the marginal.
Right: Sample average $\mathbf{\eta^{\alpha}(n)}$ changing $n_{max}$ in the unique fixed point phase. In the unique fixed point phase, if $n_{max}$ is too small, the mean abundance will show a peak in correspondence to the value $n=n_{max}-1$. Enlarging $n_{max}$ will correct the error. The mean of the distribution always remains at values near to $K=280$. Both figures refer to the same graph and interaction matrix with parameters $\hat{\mu}=0.1$, $T=0.2$, $\hat{\sigma}=0.2$, $N=200$.
In Fig. 16 we show the mean abundance on a given network and the marginal probabilities for three different species in the unique fixed point phase. If $n_{max}$ is too small, the marginals for some species will be wrong: We detect the error looking if there are some species for which $\eta_{i}(n_{max}-1)\neq 0$. However, enlarging $n_{max}$ will then correct the marginal. Once we identify a value of $n_{max}$ such that $\eta_{i}(n_{\text{max}})=0$ inside computer precision, enlarging $n_{max}$ further will have no effect on the marginals: this is a clear indication of the fact that the solution found for the BP equations is not influenced by our numerical implementation, as long as $n_{max}$ is large enough. A good practical choice for $n_{max}$ could be $n_{max}\simeq 2K$. In fact, without interactions, the marginal for each species will be centered around $n=K$: the interactions could modify slightly the mean of the marginals, but not so much: one can verify from Fig. 16 right that the mean of the distribution remains at values near to $K=280$ in this phase.
The behaviour in the unbounded growth phase is instead different. The numerical implementation of the BP equations with finite $n_{max}$ is converging even in the unbounded growth phase. However, this should not bother the reader too much: once implemented with a finite $n_{max}$, the BP equations are searching a solution for the marginal distributions generated not by the original Hamiltonian in eq. (5), but by a modified Hamiltonian in which we are manually injecting a reflecting wall at $n=n_{max}$. While in the unique fixed point phase the final marginals are independent on the value of $n_{max}$ as long as it is large enough, as explained in the precedent paragraph, in the Unbounded Growth phase some marginals will always depend on $n_{max}$, irrespectively on how large $n_{max}$ is, and in particular they will always have $\eta_{i}(n_{max}-1)\simeq 1$, as shown in the left panel of Fig. 17: this is a clear evidence that the abundance distribution for $n_{max}\to\infty$ would run off to infinity.
Figure 17: Left: Marginals $\mathbf{\eta_{i}^{\alpha}(n)}$ for three species changing $n_{max}$ in the Unbounded Growth phase. While for some species changing $n_{max}$ does not change their marginals, like species 1 in the figure, some other species have $\eta_{i}(n_{max}-1)\simeq 1$ irrespectively on the value of $n_{max}$, like species 3.
Right: Sample average $\mathbf{\eta^{\alpha}(n)}$ changing $n_{max}$ in the Unbounded Growth phase. In the Unbounded Growth phase, the mean abundance will continue to change changing in $n_{max}$. In particular, $\eta^{\alpha}(n_{max}-1)$ will have a finite weight that, for $n_{max}$ large enough, does not depend on $n_{max}$. Both figures refer to the same graph and interaction matrix with parameters $\hat{\mu}=0.1$, $T=1$, $\hat{\sigma}=0.5$, $N=200$.
In the right panel of Fig. 17, we show the mean abundance in the Unbounded Growth phase changing $n_{max}$: it is evident that $\eta^{\alpha}(n_{max}-1)$ has a finite weight that, for $n_{max}$ large enough, does not decrease anymore increasing $n_{max}$, differently with what happened in the Unique Fixed point phase. This can be chosen as a practical rule for detecting the Unbounded Growth phase: if $\eta^{\alpha}(n_{max}-1)$ is different from zero and its value does not decrease increasing $n_{max}$ the system is in the Unbounded Growth phase. We stress that, at variance with what happens in fully connected models, there are only some species whose marginals run off to infinity while, the marginals for other species remain finite and independent on $n_{max}$: this is due to the sparsity of the model that could create subsystems whose species grow indefinitely while not influencing the others that stay finite. The convergence of BP equations even in the Unbounded Growth phase allows us to count the fraction of diverging species, that is shown in Fig. 7 of the main text.
Appendix D Comparison between Belief Propagation and Langevin dynamics simulations
This Appendix shows that the exact equilibrium marginal distributions computed by BP are fully consistent with the marginal distributions obtained from SDE simulations of the Langevin dynamics. In Fig. 18 we show this agreement for 4 different values of $\hat{\sigma}$. For each value of $\hat{\sigma}$, at $\hat{\mu}=0.1$ and $T=1$, we display both the species abundance trajectories $n_{i}(t)$ (left panels) and the corresponding equilibrium marginal distributions (right panels). In particular, the marginal distributions are obtained in two ways: directly from BP (shown in solid lines) and by sampling the long-time dynamical evolution of the abundances (shown in dots). As shown in the right panels, the two methods perfectly coincide. Small differences only appear for marginal distributions with a finite probability of extinctions (the peak at $n=0$); such differences come from the different regularizations used at $n=0$: in the SDE simulations through the reflecting wall condition $n^{\text{min}}=\lambda\ll 1$, and in BP through the parameter $\epsilon$. These differences on the regularization parameters have no impact on the phenomenology, as also discussed in Appendix B and G.
The simulations in Fig. 18 are performed on a single graph realization, which is kept identical for both BP and the dynamics. We focus on five species of the ecosystem, highlighting their temporal evolution in color, and plotting with the same color their equilibrium abundance distributions, both from BP and from the dynamics. For $\hat{\sigma}=0$, where no disorder is present, the temporal evolutions differ because of the different demographic noise sequence, but their equilibrium distributions are equivalent for all species. For $\hat{\sigma}>0$ the five marginal distributions become distinct across species, but are still in perfect agreement between BP and the dynamical sampling.
Figure 18: Left: Trajectories of the species abundances $n_{i}(t)$ in time, simulated through the Langevin dynamics. The same five species are selected for all the different sets of parameters and their temporal evolution is higlighted in color. Right: Comparison between the marginal distributions computed with BP (in solid lines) and those sampled from the dynamical trajectories (as points). The species distributions from BP and from the dynamics, plotted in the same color of the species temporal evolution, perfectly coincide.
We then show an example of another result, obtained in the main text through BP, which is perfectly consistent with the simulation of the SDE dynamics. Such result is the appearance of the unbounded growth phase presented in Fig. 17. In particular, focusing on an interaction graph with N=128 species and parameters $\hat{\mu}=0.1,T=1$, we have run the dynamics for three different values of $\hat{\sigma}=0.30,0.35,0.40$. As shown in Fig. 19, at $\hat{\sigma}=0.30$ species simply fluctuate in time always keeping finite abundances. For $\hat{\sigma}=0.35$ instead, some species start to grow indefinitely already at small times, and in particular the number of species which undergo unbounded growth increases increasing $\hat{\sigma}$, as shown in the panel of Fig. 19 with $\hat{\sigma}=0.40$. This analysis confirms what shown in Fig. 17, namely that, for $\hat{\mu}=0.1$ and $T=1$, around $\hat{\sigma}\sim 0.35$ the unbounded growth phase appears.
Figure 19: Species abundance trajectories near the unbounded growth transition. For $\hat{\mu}=0.1$ and $T=1$, for $\hat{\sigma}=0.30$ no unbounded growth is present. For $\hat{\sigma}=0.35,0.40$ instead some species in the ecosystem grow indefinitely, displaying an unbounded growth. In these simulations the interaction graph is generated for an ecosystem of $N=128$ species.
Appendix E Universality and Scaling Law for the Sparse Ecosystems Dynamics